OpenStudy (tyteen4a03):

sin(x)^2 = 1 - cos(x), 0<= x <= 360 degrees Not sure how to tackle this problem.

OpenStudy (anonymous):

sin(x)^2 = 1-cos(x)^2 helps ?

OpenStudy (tyteen4a03):

The question states cos(x), not cos(x)^2.

OpenStudy (anonymous):

i change the sin(x)^2 by 1-cos(x)^2 ..

OpenStudy (anonymous):

1-cos(x)^2 = 1 - cos(x) cos(x) - cos(x)^2 = 0 cos(x)(1 - cos(x)) = 0 ..

OpenStudy (tyteen4a03):

I see. Thanks!

OpenStudy (anonymous):

yw